OpenStudy (anonymous):

Multiply and simplify by factoring. Assume that all expressions under radicals represent nonneggative numbers. (see comments for work)

OpenStudy (anonymous):

\[\sqrt{18a^{13}b}\times \sqrt{27a ^{14}b ^{12}}\]

OpenStudy (mertsj):

\[\sqrt{18a ^{13}b}\sqrt{27a ^{14}b ^{12}}\sqrt{9(2)(3)(3)(3)a ^{27}b ^{13}}=3(3)\sqrt{a ^{26}ab ^{12}b(6)}=\]

OpenStudy (mertsj):

\[9a ^{13}b ^{6}\sqrt{6ab}\]